This is coming from the University of Houston online textbook, section 3.6 #39, if that helps anyone else. I'm not sure how to get "fog" when there is a number to multiply by.
Given the functions f (x) = x2 + 2 and
g(x) = 5x − 8 , find:
(a) f (g(1)) (b) g(f (1))
(c) f (g(x)) (d) g(f (x))
(e) f ( f (1)) (f) g(g(1))
(g) f ( f (x)) (h) g(g(x))

Answer by stanbon(48535)


Given the functions f (x) = x2 + 2 and
g(x) = 5x − 8 , find:
(a) f (g(1)) = f[5*1-8] = f[-3] = (-3)^2+2 = 11
(b) g(f (1)) = g[1+2] = g[3] = 5*3-8 = 7
(c) f (g(x)) = f[5x-8]= (5x-8)^2+2 = 25x^2-80x+66
--------
I'll leave the rest for you.
--------------------------
Cheers,
Stan H.
(d) g(f (x))
(e) f ( f (1))
(f) g(g(1))
(g) f ( f (x))
(h) g(g(x))